# Facial Expression Analysis and Its Potentials in IoT Systems: A Contemporary Survey

### v1

@@ 6. Potential Applications of Facial Expression Analysis in IoT
The convergence of facial expression analysis with IoT systems unlocks new chances for intelligent, pervasive and adaptive computing services. By interpreting human emotions, IoT based applications can provide personalized, efficient, and responsive user experiences. In this section, we explore the potential of facial expression analysis in IoT by separately examining the contributions of MaE recognition and holographic MiE analysis.
@@ 6.1. MaE and Its Applications in IoT
The recent years have seen the rapidly expanding attention on the integration of emotion recognition and intelligent IoT devices. Such integration enhances real-time data analysis, minimal communication overhead, and data protection. Muhammad et al. (Muhammad and Hossain, 2021) designed an edge computing-empowered emotion recognition system that consists of IoT devices, an edge device, and a remote cloud. Pre-processing tasks (such as face detection, cropping, contrast enhancement, and resizing) are performed on IoT devices to alleviate computation and communication expenditure. Then, the edge device leverages the processed images for inference and decision-making and returns the decisions to the IoT devices. Moreover, the edge device downloads a global deep model from the cloud during off-peak hours to minimize latency. The remote cloud manages large-scale training and updates to ensure that the system remains up-to-date. The energy efficiency, lightweight design, and strong performance in experiments of the proposed architecture demonstrate the advantages of integrating emotion recognition with IoT devices.
For the privacy-preserving objective, Chen et al. (Chen et al., 2020a) introduced an IoT-based MaE recognition method. More specifically, due to the imbalanced and limited size of expression datasets, data augmentation is employed to enhance model performance. The IoT-based MaE recognition method ensures that sensitive emotional data is processed locally in order to reduce the risks of privacy leakage. Yang et al. (Yang et al., 2021) proposed a MaE recognition system for human-computer interaction that operates efficiently in real-time. By using a lightweight deep model on IoT devices to recognize emotions through AUs, the proposed system can minimize pre-processing steps and limited data transmission. Consequently, it offers low latency and reduced overhead that satisfies the stringent requirements of real-time applications. Chen et al. (Chen et al., 2022a) developed a system combining face recognition and emotion detection for use in autonomous vehicles. Leveraging IoT devices such as cameras, the system can facial emotions and identities of drivers to enhance takeover effectiveness. Optimized IoT protocols reduce the delay to just 2 milliseconds, while achieving high reliability with an F1 score of 96%.
@@
The application of MaE analysis in the IoT-based healthcare industry has demonstrated a promising potential. By integrating the MaE information with IoT technology, healthcare systems can provide real-time analysis and feedback that are crucial for improving medical services and patient care. For example, Hossain et al. (Hossain and Muhammad, 2019) developed an emotion recognition system for medical facility enhancement. The proposed approach utilized an IoT-based audio-video framework to assist healthcare providers in evaluating services. The system performs audio feature extraction and face detection on edge devices. Then, the remote cloud delivers the well-trained parameters to edge devices for efficient model inference. Such setup allows healthcare providers to gain immediate insights into patient emotions, facilitating more personalized and responsive care. Wang et al. (Wang et al., 2022d) highlights the potential of IoT-based MaE analysis in ensuring worker safety in hazardous industries. More specifically, a deep capsule method based on IoT devices was introduced in (Wang et al., 2022d) to monitor the mental state of miners in underground mining environments. The proposed system collects facial expressions and electroencephalograms to infer mental health states. By integrating the two modalities of data, the system could detect signs of stress and fatigue and thereby reduces the risk of man-made accidents.
@@
The application of IoT-based MaE analysis in flu tracking can enhance the management and control of flu spread through real-time monitoring and data analysis. For example, Rahman et al. (Rahman and Hossain, 2021) developed an edge internet of medical things framework to capture relevant signals (e.g., facial expressions and psychological states) during pandemics. By leveraging edge computing, the proposed edge internet of medical things framework can process data locally to generate immediate reports that highlight key insights. Such setup ensures low latency, enhanced security, and robust privacy protection that are critical factors in managing sensitive health information. Moreover, the proposed framework supports a variety of applications, such as sleep analysis, face mask detection, and physiological state analysis during a pandemic.
@@ 6.2. MiE and Its Applications in IoT
Although emotion can be reflected in facial expressions, it is difficult for someone to capture the emotional changes of another person in time due to the complexity of actual situations. To capture the true emotion and provide timely feedback in negotiation, Xiong et al. (Xiong et al., 2023) introduced a real-time MiE-based emotion recognition system via wearable smart glasses. By collecting and extracting features of the facial video on the front end of smart glasses, users can understand the communicator’s intended information through real-time analysis of MiEs.
@@ 7. Challenges and Future Directions
The research on facial expression analysis has sparked numerous notable studies, which resulted in significant advances in both its application and further research. Despite its current prosperity in various domains, there remain several challenges and issues that require attention in the future. Here, we provide three notable research topics and potential research directions in the blow.
7.1. Advancing Facial Expression Datasets
Current datasets used for training effective facial expression analysis systems face significant challenges. In particular, compared to MaE, the available data for MiE is limited. This is due to the subtle and transient nature of MiE, making data collection and annotation more difficult. Most existing MiE datasets are collected in controlled laboratory settings, which may limit their generalizability to real-world applications. Additionally, variations in methodologies and paradigms used for data collection further complicate efforts to standardize and expand these datasets, hindering their scalability and growth.
Annotating MiE data also presents a major challenge. The process requires certified expertise to ensure accuracy and reliability, which significantly increases the time and resource investment needed for dataset creation. While MaE datasets tend to be larger, they also face challenges, primarily due to biases in data collection. Most MaE datasets are derived from publicly available internet sources, leading to inconsistencies in data quality and potential bias.
To address these challenges, future research should focus on several key areas: 1) Unsupervised and Semi-Supervised Learning: These approaches can reduce reliance on large, annotated datasets by enabling models to learn from limited or even unlabeled data; 2) Synthetic Data Generation: Leveraging techniques such as GANs or diffusion models can augment existing datasets with realistic facial expression samples that retain key analytical features; 3) Standardized Protocols for Data Collection: Developing universal standards for collecting and annotating facial expression data across diverse settings can improve dataset consistency and interoperability. This includes protocols designed for “wild” data collection, such as using wearable devices or ubiquitous cameras. Data captured by IoT devices can provide a more diverse and naturalistic representation of facial expressions.
7.2. Adaptation to Diverse Cultural Contexts
Facial expression analysis systems often rely on the assumption of universal emotional expressions. However, cultural differences play a significant role in how emotions are expressed and interpreted. While MiEs are considered universal indicators of hidden or suppressed emotions, their occurrence may vary across cultures. On the other hand, MaEs are strongly influenced by culturally learned norms that dictate the appropriateness of expressing certain emotions in specific contexts. To effectively adapt facial expression analysis systems to diverse cultural settings, datasets must include both MiEs and MaEs from a variety of cultural contexts. To improve accuracy and relevance, models should incorporate contextual information, such as the interaction environment and cultural norms.
This cultural adaptability becomes especially important in IoT systems, as devices like smart home assistants, surveillance cameras, and wearable technologies are deployed globally. For instance, the facial expression details that smart assistants in different cultural settings need to focus on may vary significantly. To achieve cultural adaptability in IoT systems, it is imperative to collect culturally diverse datasets, including expressions from different geographic regions, age groups, and social contexts.
MiE datasets should capture subtle, transient expressions across various cultures, while MaE datasets must reflect culturally influenced emotional intensity and duration. Furthermore, pre-trained models can be fine-tuned using culturally specific datasets through transfer learning. This approach enables IoT devices to adapt to regional needs without requiring complete retraining from scratch. Additionally, IoT devices should integrate contextual understanding into their emotion recognition pipelines. For example, wearable devices like smart glasses can combine facial expression data with environmental cues such as social setting or cultural event, to provide culturally sensitive feedback or assistance.
7.3. Data Privacy and Protection
Data privacy and security are critical concerns in the integration of facial expression analysis into IoT systems. Facial expression data often consists of images or videos of individuals, which can include sensitive biometric information. The collection, transmission, and storage of such data from distributed IoT devices can introduce significant privacy risks. While some datasets are collected under strict informed consent protocols, many IoT-derived datasets remain unregulated. The reliance on wireless communication protocols in IoT devices increases vulnerability to interception and hacking. Weak encryption and unsecured channels can expose sensitive facial data during transmission. Additionally, storing facial data in centralized servers or cloud-based systems raises the risk of data breaches. A lack of a unified global privacy protection framework further exacerbates these concerns. While regulations like the European Union General Data Protection Regulation set standards for biometric data protection, the enforcement of these regulations varies by country, complicating global compliance.
Synthetic facial expression data has emerged as a promising solution to address these privacy concerns. Using GANs, synthetic data can be generated to supplement real datasets or to enhance model performance. Moreover, with advancements in generative AI, large-scale text-to-video models (e.g., Sora) now offer the ability to create realistic facial expression data, providing a potential avenue for safeguarding privacy in future applications.
Federated learning is another promising technique for enhancing privacy in facial expression analysis. Unlike traditional centralized machine learning approaches, federated learning enables multiple devices to collaboratively train models while keeping sensitive data localized and without sharing raw data. This decentralized approach ensures that facial data never leaves the user’s device, reducing the risk of unauthorized access and data leakage. By only sharing aggregated model updates with a central server, federated learning prevents the reconstruction of individual user’s data. However, this approach requires efficient algorithms that can run on resource-constrained IoT devices, so future research should focus on optimizing federated learning techniques to balance performance, privacy, and energy efficiency.
7.4. Real-Time Processing
Real-time processing is a critical requirement for facial expression analysis in IoT applications. For instance, emotion-driven smart assistants and healthcare monitoring systems depend on real-time performance to function effectively. However, achieving efficient, reliable, and scalable real-time facial expression analysis remains a significant challenge. These challenges arise from the computational complexity of facial expression analysis algorithms, the resource constraints of IoT devices, and the difficulties associated with deploying these systems in the wild.
One of the primary challenges is the high computational cost of facial expression analysis algorithms. Tasks such as feature extraction, expression classification, and temporal analysis demand substantial processing power. Deep learning models, which are commonly used for these tasks, involve millions of parameters and require significant computational resources. IoT devices, however, are often resource-constrained in terms of processing power, memory, and battery life, making it difficult to meet the real-time processing demands of these algorithms.
Latency is another critical issue. IoT systems that rely on centralized cloud servers for processing can suffer from delays due to data transmission and network bottlenecks. These delays can undermine the responsiveness of applications, particularly in time-sensitive areas such as healthcare. Ensuring low-latency communication and processing is essential for effective real-time performance.
Edge computing presents a promising solution to these challenges. By processing data locally on IoT devices rather than sending it to centralized servers, edge computing reduces latency and conserves bandwidth. However, implementing facial expression analysis at the edge requires efficient hardware accelerators such as GPUs or specialized AI chips. Additionally, techniques like model quantization, pruning, and knowledge distillation can help optimize facial analysis models to reduce computational requirements while maintaining accuracy. Efficient data transmission protocols are also crucial for real-time performance, particularly for IoT networks with limited bandwidth and strict latency requirements.
7.5. Comprehensive Analysis
Comprehensive facial expression analysis involves multi-modal approaches that extend beyond the traditional focus on facial muscle movements. Recent studies have introduced various modalities such as electromyography (Kim et al., 2022), word (Wahid et al., 2023), electroencephalogram (Wang et al., 2023c; Zhao et al., 2024) and other psychological signals (Saffaryazdi et al., 2022) can also reveal the emotions. Developing a multi-modal analysis of facial expressions could take advantage of complementary information and obtain robust features of emotions. Furthermore, several datasets combined with physiological signals have been proposed. For example, Chen et al. (Chen et al., 2019) introduced the micro-gesture dataset for hidden emotion recognition. Zou et al. (Zou et al., 2022) proposed a facial expression dataset with synchronized physiological signals. More valuable research can explore the intrinsic relationship between facial expressions and other signals in the future. In addition to applying multi-modal expression analysis methods to IoT, IoT needs to consider multi-tasks for collaborative analysis.
In the context of the IoT, multi-modal expression analysis should also address the challenges of multi-tasking. IoT devices often need to process data from diverse sensors and perform various analyses simultaneously. This necessitates the development of efficient multi-task learning models to enhance system performance and responsiveness. For example, a smart home system might simultaneously analyze data from security cameras, environmental sensors, and facial expressions to enable collaborative decision-making.
Future directions for facial expression analysis in IoT should prioritize the integration of multi-modal and multi-task capabilities. By incorporating various physiological signals, researchers can develop more accurate and comprehensive emotion recognition systems. In addition, by improving the efficiency of multi-task learning models, IoT devices can maintain high performance and responsiveness while executing complex tasks. This will enable IoT systems to better understand and respond to emotional demands of users to provide more intelligent and human-centric services.

### v3

@@ 6. Potential Applications of Facial Expression Analysis in IoT
The convergence of facial expression analysis with IoT systems unlocks new chances for intelligent, pervasive and adaptive computing services. By interpreting human emotions, IoT-based applications can provide personalized, efficient, and responsive user experiences. In this section, we explore the potential of facial expression analysis in IoT by separately examining the contributions of MaE recognition and holographic MiE analysis.
@@ 6.1. MaE and Its Applications in IoT
Recent years have witnessed the integration of emotion recognition with intelligent IoT devices for enhancing real-time performance, minimizing communication overhead, and protecting data privacy. Muhammad et al. (Muhammad and Hossain, 2021) proposed an edge-computing-based system that performs pre-processing on IoT devices and inference on edge servers. Pre-processing tasks (e.g., face detection, cropping, contrast enhancement, and resizing) can be performed on IoT devices to alleviate computation and communication expenditure. Then, the edge device leverages the processed images for inference and decision-making and returns the decisions to the IoT devices. Moreover, the edge device downloads a global deep model from the cloud during off-peak hours to minimize latency. Chen et al. (Chen et al., 2020a) ensured privacy by processing sensitive emotional data locally. Yang et al. (Yang et al., 2021) developed a real-time system using lightweight models and minimal transmission. Chen et al. (Chen et al., 2022b) applied emotion detection in vehicles, achieving 2 ms latency and a 96% F1 score through optimized IoT protocols.
@@
The application of MaE analysis in the IoT-based healthcare industry has demonstrated a promising potential. By integrating the MaE information with IoT technology, healthcare systems can provide real-time analysis and feedback that are crucial for improving medical services and patient care. For example, Hossain et al. (Hossain and Muhammad, 2019) developed an emotion recognition system for medical facility enhancement. The proposed approach utilized an IoT-based audio-video framework to assist healthcare providers in evaluating services. The system performs audio feature extraction and face detection on edge devices. Then, the remote cloud delivers the well-trained parameters to edge devices for efficient model inference. Such setup allows healthcare providers to gain immediate insights into patient emotions, facilitating more personalized and responsive care. Wang et al. (Wang et al., 2022b) highlights the potential of IoT-based MaE analysis in ensuring worker safety in hazardous industries. More specifically, a deep capsule method based on IoT devices was introduced in (Wang et al., 2022b) to monitor the mental state of miners in underground mining environments. The proposed system collects facial expressions and electroencephalograms to infer mental health states. By integrating the two modalities, the system could detect signs of stress and fatigue and thereby reduces the risk of man-made accidents.
@@
The application of IoT-based MaE analysis in flu tracking enhances pandemic management through real-time monitoring and data analysis. For instance, Rahman et al. (Rahman and Hossain, 2021) developed an edge Internet of Medical Things framework to capture signals (e.g., facial expressions and psychological states) during pandemics. By leveraging the power of edge computing, the framework processes data locally to generate immediate reports with key insights. The local processing capability ensures low latency, privacy preservation, and enhanced security that are critical for managing sensitive health data. Additionally, the framework supports various applications, e.g., sleep analysis, face mask detection, and physiological state monitoring during pandemics.
@@ 6.2. MiE and Its Applications in IoT
Although emotion can be reflected in facial expressions, it is difficult for someone to capture the emotional changes in time due to the complexity of actual situations. To capture the true emotion and provide timely feedback in negotiation, Xiong et al. (Xiong et al., 2023) introduced a real-time MiE-based emotion recognition system via wearable smart glasses. By collecting and extracting features of the facial video on the front end of smart glasses, users can understand the communicator’s intended information through real-time analysis of MiEs.
@@ 7. Challenges and Future Directions
Despite considerable progress in facial expression analysis, the path toward practical, real-world deployment remains fraught with challenges that include data-related limitations (e.g., small size, low inter-rater agreement), model-related issues (e.g., overfitting and low generalization), and system-level constraints (e.g., user privacy and real-time processing). More specifically, facial expression datasets remain limited due to their transient nature and complex annotation requirements (Muhammad and Hossain, 2021). Most MiE datasets are collected under controlled conditions, restricting real-world applicability. Future work should emphasize unsupervised learning, synthetic data generation using GANs or diffusion models, and standardized data collection protocols to improve scalability. Cultural adaptation remains crucial. While MiEs are considered cross-culturally consistent, MaEs vary widely due to cultural display rules (Yang et al., 2021). Regionally diverse datasets and contextual metadata (e.g., social setting and hierarchy) are necessary for improving model generalizability. Transfer learning and contextual fusion with environmental signals can enable systems to offer culturally appropriate responses. data privacy is a major concern in IoT-based facial analysis. Facial data, being biometric, poses security risks during collection and transmission via unsecured or wireless channels. Synthetic data generation (e.g., with Sora) and federated learning offer privacy-preserving alternatives, keeping data local while enabling collaborative training (Chen et al., 2020a). However, optimization is needed for deployment on resource-constrained devices. Real-time processing is essential in applications like emotion-aware assistants or medical monitoring. High latency from cloud processing must be avoided. Solutions include edge computing, model compression techniques (e.g., quantization, pruning, distillation), and event-driven models using neuromorphic cameras (Chen et al., 2022b). Lightweight communication protocols are vital to sustain real-time performance in bandwidth-limited scenarios. Finally, comprehensive analysis using multi-modal inputs such as electromyography (Kim et al., 2022), electroencephalogram (Wang et al., 2023b; Zhao et al., 2024), and psychological cues (Saffaryazdi et al., 2022) can enhance emotion recognition. Efficient multi-task learning models are also needed to manage diverse sensor inputs in complex IoT environments (Chen et al., 2019; Zou et al., 2022). These efforts will support human-centric, context-aware services in future IoT systems. Comprehensive discussions can be found in Appendix 5.